OpenStudy (anonymous):

Given the linear equation y = x – 7, find the y-coordinates of the points (-2, ), (7, ), and (2, ). Please show all of your work. Plot these points and graph the linear equation.

OpenStudy (anonymous):

-9,0,-5 just substitute the value of x

OpenStudy (anonymous):

y=(-9,0,-5)

OpenStudy (anonymous):

just substitute the value of x to the variable x.

OpenStudy (anonymous):

can you see me please

OpenStudy (anonymous):

can you see me please

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

can you show me how

OpenStudy (anonymous):

y=x-7 the coordinates that were given was (-2,y), (7,y) and (2,y) the coordinate point contains two no. the x and y.

OpenStudy (anonymous):

you were ask to find the value of y right.

OpenStudy (anonymous):

so all you've got ot do is substitute the values of x to the equation.

OpenStudy (anonymous):

y=-2-7 y=7-7 y=2-7 and uala! you ca now already solve for the values of y